The question is Sara takes 2 friends bread, she gives Judy 3times the aomunt then Jane. The total amount of loafs are 12. How many does judy get?
Answer by ptaylor(2198)

Let x=number of loaves of bread that Sara gives Jane
Then Sara gives Judy 3x loaves
Now we are told that the total amount of loaves is 12,
so:
x+3x=12
4x=12 divide both sides by 4
x=3 ----------------number of loaves of bread that she gives Jane
3x=3*3=9-----------------number of loaves that Judy gets
CK
9+3=12
12=12
and
9=3*3
9=9
Hope this helps---ptaylor